# The cool and the cruel: separating hard parts of LWE secrets

### v1

@@ 6 Recovery Results
Avg $4.81\times 10^{7}$ $9.63\times 10^{1}$ $6.85\times 10^{0}$ $4.08\times 10^{4}$ $4.61\times 10^{4}$ $3.85\times 10^{0}$
Note that for RLWE, the Time is multiplied by $n$ to account for the time for running the $n$ brute force attacks, although they can be run simultaneously. The final row provides the average values across all secrets. As can be seen, the RLWE attack is significantly faster on average, with the ratio of average times being $\sim 10^{3}$ for $n=256$ and $5\times 10^{3}$ for $n=512$, but converge to the theoretical values in Figure 3 when estimated on more secret samples. The RLWE advantage can also be seen in the values of $h^{(1)}_{u}(s)$ which are much smaller than those of $h_{u}(s)$.
Table 9: Attack times (in seconds) and estimated costs for $n=512,n_{u}=75,h=60,\log q=41$ for random $20$ secrets sorted by number of secret cruel bits in LWE. We use $40K$ samples for cruel bit recovery and $1.5M$ for Algorithm 1
@@
5 $2.82\times 10^{5}$ 9 $1.82\times 10^{5}$ $4.62\times 10^{4}$ 4
6 $2.82\times 10^{5}$ 9 $1.82\times 10^{5}$ $5.01\times 10^{4}$ 4
7 $2.82\times 10^{5}$ 9 $1.95\times 10^{5}$ $8.72\times 10^{4}$ 5
@@
13 $1.36\times 10^{7}$ 11 $1.82\times 10^{5}$ $6.21\times 10^{4}$ 4
14 $1.36\times 10^{7}$ 11 $1.95\times 10^{5}$ $9.58\times 10^{4}$ 5
@@
Avg $1.02\times 10^{9}$ $10.1$ $2.01\times 10^{5}$ $4.05$
@@ 8 Discussion and Future Work
We have presented an attack on Learning With Errors in the setting of sparse binary secrets. Our key insight is the reduction of the LWE problem to one of smaller dimension via lattice reduction and sufficient subsampling. We discuss the attack empirically and theoretically and highlight actual successful attacks for dimension up to $n=768$. We leave several important tasks as future work.
@@ 0.A.1 Rotating Reduced Short 2-cyclotomic RLWE Vectors
where $X$ is defined as $X=\text{Circ}(x_{\text{vec}})^{\top}$, with Circ denoting the skew-circulant and $x$ being the degree 1 polynomial whose vector representation is given by $x_{\text{vec}}=(0,1,0,\dots,0)\in\mathbb{Z}^{n}$. We further introduce the notation $\Phi_{k}(A)$ to denote the $k$-fold composition of $\Phi$, i.e., $\Phi^{k}(A)=\Phi\circ\Phi\circ\dots\circ\Phi$.
if $A_{\text{circ}}=\text{Circ}(a)$, then $A_{\text{circ}}$ and $X$ commute: $XA_{\text{circ}}=A_{\text{circ}}X$. This is due to the associativity in the ring $R_{q}$.
This allows us to have new samples $(A^{\rightarrow k},b^{\rightarrow k})$ by shift-negating the reduced samples $k$ times:
$A^{\rightarrow k}:=AX^{k}=\Phi_{k}(A)$
and
@@
We illustrate empirically how subselecting samples from different indicies of the circulant matrix affects $h_{u}$ in Figure 4. When we subsample elements when $k=0$ (e.g. no shifting), $h_{u}=16$ for a $n=256$ $\log q=12$, $h=20$ problem. However, subsampling elements at $k=121$ yields $h_{u}=4$, a much easier secret.
@@ 0.A.2 Minimum Sample Requirement for Hypothesis Testing
[…]
@@
We consider the unbiased variance estimator for a known 0 mean: $\hat{\sigma}^{2}_{M}=\frac{1}{M}\sum_{i=0}^{M-1}x_{i}^{2}$.
The threshold $\iota(\alpha,M)$ is set so that the type-1 error: $\alpha=\mathbb{P}_{\mathcal{U}_{q}}(\hat{\sigma}^{2}_{M}<\iota(\alpha,M))=\mathbb{P}(\hat{\sigma}^{2}_{M}<\iota(\alpha,M)|{x\sim\mathcal{U}_{q}})$ is given. Using Central Limit Theorem and solving for $\iota(M,\alpha)$
@@ 0.A.3 Our observation and the q-ary lattice z-shape
To avoid confusion, we comment briefly on a similar looking phenomenon. Prior work has observed a "z-shape" in q-ary lattices and leveraged it in the so-called hybrid attack [22, 3]. This classic z-shape is exhibited by the Gram-Schmidt orthogonalized rows of a q-ary lattice before and after lattice reduction (see left plot of Figure 5). This is distinct from the behavior we observe in the columns of $\mathbf{A}$ after reduction (right figure of Figure 5).
Figure 5: The log-norms of Gram-Schmidt orthogonalized rows of q-ary embedded $\mathbf{A}$ (left) and the stdev of columns of $\mathbf{A}$ after it is extracted from the q-ary embedding. Both results are for $n=256$, $\log_{2}q=12$ LWE problems, using the same $\mathbf{A}$ matrix.

### v2

@@ 6 Recovery Results
We conduct several experiments comparing LWE with RLWE, using identical secrets. The results of these experiments, for parameters $n=256$, $n=512$ and $n=1024$, are presented in Tables 8, 9, and 10. Each table row corresponds to a distinct secret $s$, and the columns detail the actual/estimated time costs for the LWE/RLWE attacks, along with the number of cruel bits in the secret. In the tables, Actual Time is the empirically measured duration of the attack, while the Estimated Time is estimated using an approximation $T_{\text{guess}}^{\text{LWE}}=b+a\sum_{k=0}^{h_{u}(s)}\binom{n_{u}}{k}$. This estimate provides an idea of the time required to attack a secret when $h_{u}$ is large.
For RLWE attacks, time is multiplied by $n$ since $n$ brute force attacks are run, although they can be run simultaneously. The final row of Table 8 provides the average values across all secrets. As can be seen, the RLWE attack is significantly faster on average, with the ratio of average times being $\sim 10^{3}$ for $n=256$ and $5\times 10^{3}$ for $n=512$, but converge to the theoretical values in Figure 3 when estimated on more secret samples. The RLWE advantage can also be seen in the values of $h^{(1)}_{u}(s)$ which are much smaller than $h_{u}(s)$.
Table 9: Attack times vs. estimated costs for $n=512,n_{u}=75,h=60,\log q=41$ (in seconds) for $10$ random secrets sorted by ##\## cruel bits in the secret. We use $40K$ samples for cruel bit recovery and $1.5M$ for Algorithm 1
@@
Table 10: Attack times vs. estimated times for $n=1024,h=20,\log q=50,n_{u}=495$ (in seconds) for $10$ random secrets sorted by ##\## cruel bits in the secret. We use $10K$ samples for cruel bit recovery (scaled to 4 GPUs) and $100K$ for Algorithm 1
LWE RLWE
Secret Estimated Actual Estimated Actual
time (sec) time (sec) $h_{u}(s)$ time time $h^{(1)}_{u}(s)$
1 $6.34\times 10^{9}$ 8 $1.51\times 10^{9}$ 6
2 $6.34\times 10^{9}$ 8 $2.01\times 10^{7}$ $2.01\times 10^{7}$ 5
3 $3.45\times 10^{11}$ 9 $1.71\times 10^{6}$ $4.78\times 10^{5}$ 3
4 $3.45\times 10^{11}$ 9 $1.51\times 10^{9}$ 6
5 $3.45\times 10^{11}$ 9 $1.06\times 10^{11}$ 7
6 $1.68\times 10^{13}$ 10 $1.06\times 10^{11}$ 7
7 $1.68\times 10^{13}$ 10 $1.51\times 10^{9}$ 6
8 $1.68\times 10^{13}$ 10 $1.06\times 10^{11}$ 7
9 $3.02\times 10^{16}$ 12 $1.51\times 10^{9}$ 6
10 $1.13\times 10^{18}$ 13 $1.90\times 10^{6}$ $3.14\times 10^{6}$ 4
@@ 8 Discussion and Future Work
We have presented an attack on Learning With Errors in the setting of sparse binary secrets. Our key insight is the reduction of the LWE problem to one of smaller dimension via lattice reduction and sufficient subsampling. We discuss the attack empirically and theoretically and highlight actual successful attacks for dimension up to $n=1024$. We leave several important tasks as future work.
@@ 0.A.1 Rotating Reduced Short 2-cyclotomic RLWE Vectors
where $X$ is defined as $X=\text{Circ}(x_{\text{vec}})^{\top}$, with Circ denoting the skew-circulant and $x$ being the degree 1 polynomial whose vector representation is given by $x_{\text{vec}}=(0,1,0,\dots,0)\in\mathbb{Z}^{n}$. We further introduce the notation $\Phi_{k}(A)$ to denote the $k$-fold composition of $\Phi$, i.e., $\Phi^{k}(A)=\Phi\circ\Phi\circ\dots\circ\Phi$. If $A_{\text{circ}}=\text{Circ}(a)$, then $A_{\text{circ}}$ and $X$ commute: $XA_{\text{circ}}=A_{\text{circ}}X$. This is due to the associativity in the ring $R_{q}$. This allows us to have new samples $(A^{\rightarrow k},b^{\rightarrow k})$ by shift-negating the reduced samples $k$ times:
$A^{\rightarrow k}:=AX^{k}=\Phi_{k}(A)\quad\text{ and}$
@@
We illustrate empirically how sub-selecting samples from different indices of the circulant matrix affects $h_{u}$ in Figure 4. When we sub-sample elements when $k=0$ (e.g. no shifting), $h_{u}=16$ for a $n=256$ $\log q=12$, $h=16$ problem. However, sub-sampling elements at $k=124$ yields $h_{u}=4$, a much easier secret.
@@ 0.A.2 Minimum Sample Requirement for Hypothesis Testing
We examine two centered discrete distributions, namely $\mathcal{U}_{q}:=\textit{Uniform}(\mathbb{Z}_{q})$ and $\mathcal{G}_{q}:=\textit{Discrete Gaussian}(0,\sigma_{\mathcal{G}}^{2})\mod q$ where $\sigma_{\mathcal{G}}^{2}=h_{r}\sigma_{r}^{2}+\sigma_{e}^{2}$ and $\sigma_{\mathcal{G}_{q}}^{2}=F_{q}(\sigma_{\mathcal{G}}^{2})$.
@@
We consider the unbiased variance estimator for a known 0 mean: $\hat{\sigma}^{2}_{M}=\frac{1}{M}\sum_{i=0}^{M-1}x_{i}^{2}$. The threshold $\iota(\alpha,M)$ is set so that the type-1 error: $\alpha=\mathbb{P}_{\mathcal{U}_{q}}(\hat{\sigma}^{2}_{M}<\iota(\alpha,M))=\mathbb{P}(\hat{\sigma}^{2}_{M}<\iota(\alpha,M)|{x\sim\mathcal{U}_{q}})$ is given. Using Central Limit Theorem and solving for $\iota(M,\alpha)$
@@ 0.A.3 Our Observation and the q-ary Lattice Z-shape
Prior work has observed a "z-shape" in q-ary lattices in the context of the hybrid attack [24, 4]. This classic z-shape is exhibited by the Gram-Schmidt orthogonalized rows of a q-ary lattice (see left plot of Figure 5, orange before reduction, blue after). This is distinct from the behavior we observe in the columns of $\mathbf{A}$ after reduction (see right plot of Figure 5 in blue).
Figure 5: